A monopolistic producer of two goods, G1 and G2, has a total cost function

TC = 5Q(1) +10Q(2)

where Q(1) and Q(2) denote the quantities of G1 and G2 respectively. If P1 and P2 denote the corresponding prices then the demand equations are


P(1) = 50 - Q(1) - Q(2)

P(2) = 100 - Q(1) - 4Q(2)


Find the maximum profit if the firm's total costs are fixed at $100. Estimate the new optimal profit if total costs rise to $101.


1
Expert's answer
2021-11-29T19:10:03-0500

The profit-maximizing quantity is produced at MR = MC.

"MR1 = TR'(Q1) = 50 - 2Q1 - Q2,"

"MC1 = TC'(Q1) = 5,"

50 - 2Q1 - Q2 = 5,

2Q1 + Q2 = 45.

"MR2 = TR'(Q2) = 100 - Q1 - 8Q2,"

"MC2 = TC'(Q2) = 10,"

100 - Q1 - 8Q2 = 10,

Q1 + 8Q2 = 90,

Q1 = 90 - 8Q2.

If we substitute the above equation into the previous one, then:

"2(90 - 8Q2) + Q2 = 45,"

15Q2 = 135,

Q2 = 9 units,

Q1 = 90 - 8×9 = 18 units.

P1 = 50 - 18 - 9 = 23,

P2 = 100 - 18 - 4×9 = 46.

Total profit is:

"TP = TR - TC = (23\u00d718 + 46\u00d79) - (5\u00d718 + 10\u00d79) = 648."

If FC rise to $101, then the total profit will decrease.